Combine the nullifier and parameter variation methods to solve the
PVI:

3y'' − 6y' + 30y = 15 sin(x) + e^x tan(3x)

y(0) = 0

y' (0) = 1
Answer #1

Solution :- Given that, = isns en ca) te^r tan (32) 39 - by +304 With Y(0)=0 yCO) = 1 yl - 2y + 10g = 5 sen att extan 3x AQ S y po - A casa + B e8qa IP s A sina t Beosu ypa - A casa -B sina 19 1990 - 4 cosu-3889 2-2 (cAsempet Bedok +10 (n casa + Bw . From 2 cosa t 3 17. 17 Gp, sina Part- The prênciple of Varlation of parcayeters to find the particulan election the deffeof parameterp. By method of (9 of vare ate Partöezlar solutions Y Pz - A ex casa t BR? Lena 9 where, 4-14 82 flse) WC) da ) BFrom aq? Y pz fin (reen Itana) + siga] e cosu to com .essence 42-gen (seenet fann Jocos ne particrelare solection of from eco yeos=o a) C1 + 2 2 = 0 -) 1 = 77 44 y!60)= 5) 161-0)+(2(1+0) 70+47 -0-0-$(1= 1 - Citlat q 17 3. -- Citle = 1 - 1 / 2 + 1 -